OpenStudy (anonymous):

evaluate integration from 0 to infinity e^-x^2

OpenStudy (anonymous):

looks like error function(x)

OpenStudy (anonymous):

\[\int\limits_{0}^{infinity} e ^{x ^{2}}\]

OpenStudy (anonymous):

hold on its -x

OpenStudy (anonymous):

\[\int\limits_{0}^{infinity} e ^{-x ^{2}}\]

OpenStudy (anonymous):

erf(inf)=1

OpenStudy (anonymous):

use the substitution x^2=u and use integratoin by parts

OpenStudy (anonymous):

it is to be done by gamma function it says

OpenStudy (anonymous):

no way you are going to get a nice closed form for this

OpenStudy (anonymous):

help me evaluate

OpenStudy (anonymous):

ok.. whenyou do the substitution i said, you'd get a gamma function \[ x^2=t\quad dx={dt \over 2\sqrt{t}}={t^{-1/2}dt\over2}\\ I={1\over2}\int_0^\infty t^{-1/2}e^{-t}dt={\Gamma(1/2)\over2} \]

OpenStudy (anonymous):

i think you have to go to another dimension i.e. introduce \(y\) honestly i forget, but you could google the proof, the answer is well known \(\frac{\sqrt{\pi}}{2}\)

OpenStudy (anonymous):

and \[\Gamma(1/2)=\sqrt{\pi}\]

OpenStudy (anonymous):

got it electro right!!!!

OpenStudy (anonymous):

wow much snappier!!! i have only seen it done this way http://www.youtube.com/watch?v=fWOGfzC3IeY

OpenStudy (anonymous):

that is from -infinity to + infinity

OpenStudy (anonymous):

it is even, take half now all we need is the proof that \(\Gamma(\frac{1}{2})=\sqrt{\pi}\)

OpenStudy (anonymous):

well i figured it out x) bt thanks